# FreeTacMan: Robot-free Visuo-Tactile Data Collection System for Contact-rich Manipulation

### v1

@@ Abstract
Enabling robots with contact-rich manipulation remains a pivotal challenge in robot learning, which is substantially hindered by the data collection gap, including its inefficiency and limited sensor setup. While prior work has explored handheld paradigms, their rod-based mechanical structures remain rigid and unintuitive, providing limited tactile feedback and posing challenges for human operators. Motivated by the dexterity and force feedback of human motion, we propose FreeTacMan, a human-centric and robot-free data collection system for accurate and efficient robot manipulation. Concretely, we design a wearable data collection device with dual visuo-tactile grippers, which can be worn by human fingers for intuitive and natural control. A high-precision optical tracking system is introduced to capture end-effector poses while synchronizing visual and tactile feedback simultaneously. FreeTacMan achieves multiple improvements in data collection performance compared to prior works, and enables effective policy learning for contact-rich manipulation tasks with the help of the visuo-tactile information. We will release the work to facilitate reproducibility and accelerate research in visuo-tactile manipulation.
Figure 1: Overview of FreeTacMan. FreeTacMan is a robot-free, human-centric visuo-tactile data collection system, featuring low-cost, high-resolution tactile sensors and a portable, cross-embodiment modular design. FreeTacMan transfers human visual perception, tactile sensing, and motion control skills to robots efficiently by integrating visual and tactile data. We showcase several contact-rich manipulation tasks that are challenging for vision-only approaches.
1 Introduction
@@
Humans inherently rely on the integration of vision and touch to perform contact-rich manipulation tasks. While vision provides comprehensive object recognition and pose estimation capabilities, tactile feedback conveys critical information about local contacts that cannot be obtained visually, such as surface texture [1, 2, 3], in-hand pose [4, 5, 6], material compliance [7, 8], and force distribution [9, 10]. For example, when handling a fragile or deformable object, vision guides initial motion planning and object localization, while tactile information enables modulation of grip force and adjustment of object orientation to prevent damage.
@@
The vision-based imitation learning has shown strong potential in robot manipulation tasks [11, 12, 13, 14], benefiting from the growing large-scale demonstration datasets [15, 16]. In the tactile domain, the lack of high-quality tactile datasets and compatible sensing hardware has prevented similar advances. Early efforts concentrated on the effectiveness of tactile information for fine-grained and contact-rich manipulation skills, with tactile data collected by sensors attached to robots [17, 18, 19, 20]. However, this data collection scheme limits the flexibility for scaling visuo-tactile data. To accelerate research in this domain, two prerequisites would be addressed: 1) A visuo-tactile sensor that can be fabricated and deployed at scale, and 2) a visuo-tactile data collection system that provides real-time tactile feedback and enables rapid redeployment.
@@
As to visuo-motor sensors, existing optical tactile sensors [21, 22, 23] are often bulky, limiting their integration with diverse end-effectors, given the cost to fabricate and narrow Field of View (FoV). Moreover, they lack the modularity required for rapid adaptation across diverse embodiments. On the data collection side, some prior protocols are not optimized for high-fidelity tactile data collection. Systems that rely on motion-capture with AR/VR visualization [24, 25, 26] or primary-replica teleoperation rigs [27, 28, 29] capture accurate camera views and joint trajectories. They conversely offer no direct, real-time tactile signals, and impose fixed robot setup with complex calibration or high latency. Handheld data collection paradigms [30, 31] release the human operators from robot embodiment, and yet transmit tactile cues through long mechanical linkages, hindering direct tactile feedback and precise measurement of instantaneous tactile signals.
@@
In this work, we introduce FreeTacMan, a robot-free and human-centric visuo-tactile data collection system to acquire robot manipulation data accurately and efficiently. As the basis of the visuo-tactile sensing, we develop novel internal optics and gel interface of the sensor to increase spatial resolution and FoV. To support efficient adaptation across end-effectors, the sensing module is decoupled from its mounting bracket via an electrical and mechanical connector. For the visuo-tactile data collection system that attaches the sensor, an in-situ111 “In-situ” indicates that FreeTacMan is designed to preserve the natural interaction between the human fingertip and the environment. It emphasizes maintaining the finger’s original grasping posture while simultaneously capturing tactile feedback. This is akin to the concept in materials science [32] and biology [33], where systems are studied in their native conditions—without altering their environment or state—to gain a deeper understanding of their behavior under actual working conditions. hardware structure is designed to guarantee precise, real-time tactile feedback during demonstration via two mechanisms - a gripper-finger interface coupled to human fingertips, and a linear transmission mechanism enabling precise gripper control and unattenuated tactile feedback. To validate the quality of data collected by FreeTacMan, we train and deploy imitation learning policies on challenging contact-rich manipulation tasks. Furthermore, the effectiveness of a temporal-aware tactile pretraining strategy is validated as well.
@@
In summary, our main contributions are: (i) A portable, high-resolution, low-cost visuo-tactile hardware sensor designed for rapid adaptation across multiple robotic end-effectors. (ii) An in-situ, robot-free, real-time tactile data-collection system that leverages a handheld end effector and the proposed sensor to excel at diverse contact-rich tasks efficiently. (iii) Experimental validation shows that imitation policies trained with our visuo-tactile data achieve an average 50% higher success rate than vision-only approaches in a wide spectrum of contact-rich manipulation tasks.
@@
Table 1: Comparison with existing data collection systems. The wearable, in-situ design of FreeTacMan delivers direct and high-precision tactile feedback to human operators. To evaluate tactile feedback fidelity of handheld methods quantitatively, we measure the number of mechanical transmission “link” inside the handheld gripper from human hand to the grasped object.
@@
Teleop: VR/AR ARCap [24] VR Controller
@@
Teleop: Primary–Replica ALOHA [28] Puppet Arm
@@
Handheld UMI [30] Trigger Contact (4 links)
@@
2 Related Work
@@
Dataset Collection System for Robot Learning. Recent robot learning has been greatly advanced through imitating expert demonstrations. Popular approaches [11, 12, 36, 13] leverage large-scale visuomotor datasets (pairs of RGBD sensors and actions) to train end-to-end policies that generalize across objects and scenes. To gather data suitable for imitation learning, prior works have relied on different interfaces to teleoperate real robots, including motion capture with AR or VR visualization [24, 16, 26, 25, 34], 3D space mouse [12, 37], primary-replica system [28, 27, 29, 35], and wearable devices [38, 39, 40]. However, these setups are either expensive, operationally complex, or suffer from limited precision and require real robot control. Handheld data collection paradigms [30, 31] offer greater freedom across robot embodiment and gain access to in-the-wild data collection. Meanwhile, their trigger-based grippers and multi-link transmission designs introduce backlash between links, which blurs tactile cues and leads to inaccurate relay of gripper contact to the operator. Different from previous work, FreeTacMan implements an in-situ data collection system that enables human operators to feel gripper contact directly in real time. The comparison of control methods and approaches of tactile feedback between ours and prior work is depicted in Table 1.
Tactile Sensing for Precise Manipulation. Tactile sensing as a compensatory modality besides visual information is crucial for robot precise manipulation tasks, especially for some contact-rich manipulation tasks [41, 42, 43, 44]. Prior work on robotic tactile sensing has used force or torque sensors [45, 26, 46, 47], which are sensitive to noise and capture limited tactile information in real contact. Electrical tactile sensors can produce more realistic contact signals [48, 49, 50, 51, 52], but often suffer from low spatial resolution, crosstalk, complex fabrication, and environment interference. Camera-based visuo-tactile sensors, such as GelSight [21, 22], GelSim [23, 53], and MCTac [54], overcome these limitations by capturing the 3D deformation of a soft gel layer, achieving high resolution and capturing contact forces and surface texture. Yet, they are manufactured as one-off modules without regard for efficient adaptation onto robot end-effectors. In contrast, our visuo-tactile sensor is designed with modular interfaces both at the data collection end and the robot end-effectors, enabling rapid reconfiguration while delivering high-resolution visuo-tactile information.
3 Visuo-Tactile Data Collection System
3.1 Hardware Design
@@
Figure 2: The hardware system of FreeTacMan. Left: The in-situ gripper in the collection and execution interface respectively, with identical visual and tactile observations. Right: (a) Composition of the visuo-tactile sensor. (b) Exploded view of FreeTacMan. (c) The modular design allows for an agile switch between the collection and execution interface. (d) Human-machine interface design.
@@
Design Criteria. To address the limitations of existing data collection systems in capturing tactile information and maintaining high-fidelity tactile feedback, as well as the issues of bulky size and limited FoV in current visuo-tactile sensors, we define the following design criteria to ensure efficient collection of high-quality visuo-tactile manipulation data with intuitive feedback. (a) High-quality visuo-tactile sensing: The designed sensor must achieve competitive sensing performance while minimizing thickness and maximizing the FOV to integration flexibility and enhance spatial perception. (b) Efficient data collection: The system should minimize the tactile transmission path from human fingers to the grasped object for real-time and precise tactile feedback, while ensuring stable and fingertip-level dexterous control. (c) Scalability: A modular design architecture should be adopted to ensure compatibility across robot embodiments. (d) Usability: The system should accommodate a wide range of fingertip sizes, providing ergonomic comfort during operation.
Visuo-Tactile Sensor. As shown in Fig. 2(a), the visuo-tactile sensor integrates the sensing module-comprising circuit and electronic modules, a compact camera, programmable illumination, acrylic, and deformable gel layer-into a single unit, collectively enabling high-fidelity tactile imaging. This module snaps onto the gripper-finger interface via a mechanical connector. Compared to the widely used GelSight Mini [22], our design offers desirable merits in many dimensions, as summarized in Table 2, leading to improved adaptability to diverse robot end-effectors.
@@
In-situ Gripper. FreeTacMan achieves hinge-free operation through visuo-tactile sensors mounted on the operator’s fingertip, where the sensor layer forms the interface between skin and manipulated objects, eliminating intermediate linkages to provide zero mechanical attenuation and natural proprioception. To ensure that the unattenuated tactile feedback directly translates to operator control precision, as illustrated in Fig. 2(b), the system incorporates a linear transmission mechanism with chrome-plated steel shafts and linear bearings, constraining movement to highly accurate linear trajectories (axial deviation $\geq$ 0.02 mm). Additionally, an inverted crank-slider mechanism converts finger-driven motion into synchronized linear output, while dual parallel shafts and rolling bearings in linking joints minimize friction and lateral torque, achieving over 90% transmission efficiency.
@@
Modular Architecture. FreeTacMan system is built as three plug-and-play modules, each optimized for rapid setup and cross-embodiment compatibility: a sensor perception module for tactile data collection, a universal gripper interface (Fig. 2(c)) for robot compatibility which supports rapid assembly and disassembly of the sensor via a compact mounting mechanism, and a camera mounting scaffold to ensure stable, aligned visual feedback from wrist camera. Customized interfaces are provided for different robot arms, including a 6-DOF Piper arm for low-load tasks and 7-DOF Franka arm for high-precision, heavy-load applications. For 3D models of end-effectors, their integration on different robot arms (e.g., Piper, Franka) and full assembly renders, please visit our demo page.
Table 2: Comparison between our visuo-tactile sensor and GelSight. The proposed hardware exhibits several desirable properties, including a higher resolution, FoV, and smaller size in terms of thickness. Though not strictly compared, our design offers a significant cost advantage over the GelSight counterpart (market price).
Hardware Resolution Ratio FoV Thickness Cost
(microns) $\downarrow$ (mm2) $\uparrow$ (mm) $\downarrow$ (USD) $\downarrow$
GelSight Mini [22] 30 – 100 265.98 28 504
Ours 15 – 20 358.47 18 30
@@
Human-Machine Interface Design. To achieve rapid adaptability, we incorporate hook-and-loop straps as a fingertip fastening mechanism, as illustrated in Fig. 2(d). These straps accommodate hand sizes ranging from the 5% to 95% percentile of adults and support repeated use [55], balancing operational efficiency. Furthermore, nylon straps offer enhanced breathability and sweat-resistant properties, ensuring user comfort and minimizing the risk of skin irritation during prolonged usage.
3.2 Human-to-Robot Data Transfer
@@
To facilitate human-to-robot skill transfer, a high-precision OptiTrack [56] motion capture system is used for 6D pose tracking of the interface at 300 Hz, achieving a mean error of 0.118 mm. Five retro-reflective markers are mounted on the interfaces, with three positioned on the top plate to measure pose and two on the grippers to capture relative displacement. The coordinates of markers are first transformed from the world coordinate frame to the robot base coordinate frame. A local coordinate frame, aligned with the robot URDF, is established at the gripper’s Tool Center Point (TCP) using three top-plate markers, allowing derivation of end-effector pose while ensuring consistency with the robot kinematic model. We downsample the tracking data to synchronize with RGB images. To this end, each frame contains: the wrist camera RGB image, two visuo-tactile images, the end-effector pose of the in-situ gripper in the world coordinate frame and gripper width. Each trajectory consists of a sequence of such embodiment-agnostic data synchronized at 30 Hz.
@@
Unlike prior works  [30, 31] relying on SLAM and IMU fusion to estimate absolute and relative end-effector poses, the motion capture system we have avoids IMU drift and tracking errors. Given the URDF of a target embodiment, we could utilize IKPY [57] as an inverse kinematic solver to map poses of the in-situ gripper to absolute joint positions directly as the action representation.
3.3 Tactile Pretraining and Policy Learning

### v2

@@ Abstract
Enabling robots with contact-rich manipulation remains a pivotal challenge in robot learning, which is substantially hindered by the data collection gap, including its inefficiency and limited sensor setup. While prior work has explored handheld paradigms, their rod-based mechanical structures remain rigid and unintuitive, providing limited tactile feedback and posing challenges for human operators. Motivated by the dexterity and force feedback of human motion, we propose FreeTacMan, a human-centric and robot-free data collection system for accurate and efficient robot manipulation. Concretely, we design a wearable gripper with dual visuo-tactile sensors for data collection, which can be worn by human fingers for intuitive control. A high-precision optical tracking system is introduced to capture end-effector poses while synchronizing visual and tactile feedback simultaneously. We leverage FreeTacMan to collect a large-scale multimodal dataset, comprising over 3000k paired visual-tactile images with end-effector poses, 10k demonstration trajectories across 50 diverse contact-rich manipulation tasks. FreeTacMan achieves multiple improvements in data collection performance compared to prior works, and enables effective policy learning for contact-rich manipulation tasks with self-collected dataset. The full suite of hardware specifications and the dataset will be released to facilitate reproducibility and support research in visuo-tactile manipulation.
I INTRODUCTION
@@
Humans inherently rely on the integration of vision and touch to perform contact-rich manipulation tasks. While vision provides comprehensive object recognition and pose estimation capabilities, tactile feedback conveys critical information about local contacts that cannot be obtained visually, such as surface texture [1], in-hand pose [2], material compliance [3], and force distribution [4]. For example, when handling a fragile or deformable object, vision guides initial motion planning and object localization, while tactile information enables modulation of grip force and adjustment of object orientation to prevent damage.
@@
The vision-based imitation learning has shown strong potential in robot manipulation tasks [5, 6], benefiting from the growing large-scale demonstration datasets [7, 8]. In tactile domain, visuo-tactile sensors offer high-resolution, sensitive, and easily integrable multimodal signals, making them particularly well-suited for existing policy learning pipelines. However, the lack of large-scale, high-quality tactile datasets and compatible sensing hardware has prevented similar advances. To accelerate research in this domain, two prerequisites would be addressed: 1) a visuo-tactile data collection system that provides real-time tactile feedback and enables rapid redeployment, and 2) a large-scale, high-precision visuo-tactile dataset covering diverse contact-rich manipulation tasks.
@@
Early efforts concentrated on collecting data using sensors mounted on robots [9, 10, 11], which limit the flexibility for scaling visuo-tactile data. Systems that rely on motion-capture with AR/VR visualization [12, 13] or primary-replica teleoperation rigs [14] capture accurate camera views and trajectories. They conversely offer no direct, real-time tactile signals, and impose fixed robot setup with complex calibration or high latency. Handheld data collection paradigms [15, 16] release the human operators from robot embodiment. However, these methods typically rely on SLAM and IMU fusion for localization, which introduces significant positioning errors that are particularly detrimental to tactile perception. Moreover, they transmit tactile cues through long mechanical linkages, hindering direct tactile feedback and precise measurement of instantaneous tactile signals.
The inefficiency and lack of real-time tactile feedback in current data collection setups compromise data quality, limit scalability, and pose a risk of sensor damage during operation. Consequently, existing visuo-tactile datasets[17, 18, 19] are largely confined to specific perception tasks or lack the diversity and precision required for generalizable visuo-tactile policy learning across diverse manipulation scenarios.
@@
In this work, we introduce FreeTacMan, a robot-free and human-centric visuo-tactile data collection system to acquire manipulation data accurately and efficiently. With a modular sensor and in-situ111“In-situ” indicates that FreeTacMan preserves natural fingertip–environment interaction. It emphasizes maintaining the original grasp posture while capturing tactile feedback. This is akin to the concept in materials science [20] and biology [21], where systems are studied in their native conditions—without altering their environment or state—to gain a deeper understanding of their behavior under actual working conditions. hardware, it ensures precise, real-time tactile feedback via two mechanisms - a gripper-finger interface coupled to human fingertips, and a linear transmission mechanism enabling precise gripper control and unattenuated tactile feedback. This foundation enables a comprehensive visuo-tactile manipulation dataset with high precision, diverse task scenarios, and rich contact dynamics, recorded with synchronized high-resolution visual, visuo-tactile, and pose data. This extensive collection provides a robust foundation for training and evaluating generalizable visuo-tactile policies. To validate the effectiveness of the data collected by FreeTacMan, we train and deploy imitation learning policies that leverage a temporal-aware tactile pretraining strategy on challenging manipulation tasks.
In summary, our main contributions are:
(i) An in-situ, robot-free, real-time tactile data-collection system that leverages a handheld gripper with modular visuo-tactile sensors to excel at diverse contact-rich tasks efficiently.
(ii) A large-scale, high-precision (sub-millimeter) visuo-tactile manipulation dataset with over 3000k visuo-tactile image pairs, more than 10k trajectories across 50 tasks.
@@
(iii) Experimental validation shows that imitation policies trained with our visuo-tactile data achieve an average 50% higher success rate than vision-only approaches across challenging contact-rich manipulation tasks.
@@
TABLE I: Comparison with existing data collection systems. Our in-situ design delivers direct and high-precision tactile feedback to operators. To evaluate tactile feedback fidelity of handheld methods quantitatively, we measure the number of mechanical transmission “link” inside the handheld gripper from human hand to the grasped object.
@@
Teleop:
VR/AR
ARCap [12] VR Controller
@@
Teleop:
Primary–
Replica
ALOHA [14] Puppet Arm
@@
Handheld
UMI [15]
Trigger
Contact (4 links)
@@
II Related Work
II-A Dataset Collection System for Robot Learning.
@@ II Related Work
Recent robot learning has been greatly advanced through imitating expert demonstrations. Popular approaches [26, 27, 28] leverage large-scale visuomotor datasets (pairs of RGBD sensors and actions) to train end-to-end policies that generalize across objects and scenes. To gather data suitable for imitation learning, prior works have relied on different interfaces to teleoperate real robots, including motion capture with AR or VR visualization [12, 8, 13], 3D space mouse [27], primary-replica system [14, 24, 25], and wearable devices [29]. However, these setups are either expensive, operationally complex, or suffer from limited precision. Handheld data collection paradigms [15] offer greater flexibility for various robot embodiments, but their trigger-based grippers and multi-link transmission designs introduce backlash between links, which blurs tactile cues and leads to inaccurate relay of gripper contact to the operator. In contrast, FreeTacMan implements an in-situ data collection system that enables operators to feel gripper contact directly in real time. The comparison of control methods and tactile feedback between ours and prior work is depicted in Table˜I.
II-B Tactile Dataset for Contact-rich Manipulation.
Most existing visuo-tactile datasets focus on perception tasks such as 6DoF pose tracking[30], cross-modal generation[17], representation learning[19], garment-feature spatially-aligned perception[31], and tactile-semantic description[32], with limited coverage of the full manipulation process. The ObjectFolder benchmark [33] includes 4 visuo-tactile manipulation tasks, exhibiting limited scale and diversity. Moreover, most tactile manipulation datasets are collected with tactile array sensors based on piezoresistive or Hall-effect principles [34, 35, 36, 37]. Although policies trained on these dataset show promising results, the intrinsic limitations of these sensors, such as low spatial resolution, crosstalk, complex fabrication, and environment interference, can constrain policy performance and reduce deployment efficiently in real-world settings. Taken together, these limitations expose an urgent demand for a visuo-tactile dataset that offers rich, high-resolution, and scalable data spanning the entire manipulation process.
III Method
III-A Hardware Design
@@
Figure 2: Hardware system. Left: The in-situ gripper in the collection and execution interface respectively, with identical visual and tactile observations. Right: (a) Composition of the sensor. (b) Exploded view of FreeTacMan. (c) The modular design allows for an agile switch between the collection and execution interface. (d) Human-machine interface design.
@@
Design criteria. To enable efficient collection of high-fidelity tactile data, we define the following criteria. (a) Multimodal data acquisition: The system must provide competitive visuo-tactile sensing with exceptional consistency, coupled with high-precision pose tracking. (b) Efficiency: The system should minimize the tactile transmission path from human fingers to the grasped object for real-time and precise tactile feedback, while ensuring stable and fingertip-level control. (c) Scalability: A modular design architecture should be adopted to ensure compatibility across robot embodiments. (d) Usability: The system should accommodate a wide range of fingertip sizes, providing ergonomic comfort during operation.
Multimodal sensing. As shown in Figure˜2(a), we employ visuo-tactile sensors and a wrist-mounted camera to capture tactile and visual information. End-effector poses are tracked using a high-precision motion capture system, achieving sub-millimeter accuracy—an essential factor for contact-rich tasks, where even minute pose errors can result in disproportionately large deviations in tactile feedback.
@@
In-situ gripper. FreeTacMan achieves hinge-free operation through visuo-tactile sensors mounted on the operator’s fingertip, where the sensor layer forms the interface between skin and manipulated objects, eliminating intermediate linkages to provide zero mechanical attenuation and natural proprioception. To ensure that the unattenuated tactile feedback directly translates to operator control precision, as illustrated in Fig.˜2(b) , the system incorporates a linear transmission mechanism with chrome-plated steel shafts and linear bearings, constraining movement to highly accurate linear trajectories (axial deviation $\geq$ 0.02 mm). Additionally, an inverted crank-slider mechanism converts finger-driven motion into synchronized linear output, while dual parallel shafts and rolling bearings in linking joints minimize friction and lateral torque, achieving over 90% transmission efficiency.
@@
Modular architecture. FreeTacMan system is built as three plug-and-play modules, each optimized for rapid setup and cross-embodiment compatibility: a sensor perception module for tactile data collection, a universal gripper interface (Fig.˜2(c)) for robot compatibility , and a camera mounting scaffold to ensure aligned visual feedback from wrist camera. The sensor is based on the McTac design [38] and features enhanced modularity and consistency. The improved universal mechanical interface allows the same sensor unit to be used interchangeably on a data collection setup or a robotic arm. Automated fabrication ensures consistent quality. Customized interfaces are provided for different robot arms, including a 6-DOF Piper arm for low-load tasks and 7-DOF Franka arm for high-precision, heavy-load applications. For 3D models of end-effectors and their integration on different robot arms (e.g., Piper, Franka), please visit our demo page.
@@
Human-machine interface design. To achieve rapid adaptability, we incorporate hook-and-loop straps for fingertip fastening, as illustrated in Fig.˜2(d). These straps fit hand sizes ranging from the 5% to 95% percentile of adults and support repeated use [39], balancing operational efficiency. Furthermore, the system features a lightweight ($157.5,g$) and compact ($145\times 85\times 106,\text{mm}^{3}$) design that ensures comfort.
III-B Human-to-Robot Data Transfer
@@
To facilitate human-to-robot skill transfer, a high-precision motion capture system [40] is used for 6D pose tracking of the interface at 300 Hz, achieving a mean error of 0.118 mm. Five retro-reflective markers are mounted on the interfaces, with three positioned on the top plate to measure pose and two on the grippers to capture relative displacement. The coordinates of markers are first transformed from the world coordinate frame to the robot base coordinate frame. A local coordinate frame, aligned with the robot URDF, is established at the gripper’s Tool Center Point (TCP) using three top-plate markers, allowing derivation of end-effector pose while ensuring consistency with the robot kinematic model. We downsample the tracking data to synchronize with RGB images. To this end, each frame contains: the wrist camera RGB image, two visuo-tactile images, the end-effector pose of the in-situ gripper in the world coordinate frame and gripper width. Each trajectory consists of a sequence of such embodiment-agnostic data synchronized at 30 Hz.
@@
Unlike prior works  [15, 16] relying on SLAM and IMU fusion to estimate end-effector poses, the motion capture system avoids IMU drift and tracking errors. Given the URDF of a target embodiment, we could utilize IKPY [41] as an inverse kinematic solver to map poses of the in-situ gripper to joint positions directly as the action representation.
III-C Tactile Pretraining and Policy Learning